3. Cournot Competition (26 points) Consider a Cournot model. The market demand is p=130-41-42. Firm ls marginal cost is 10.
Answer #1

Page 1 3. P = 130- 47 , , C = <, = 0 A, For Firm 1: Profit : ,= (-6,2, = (130-4,-23-1022, - (120 – ,-22)9, = (120-92), - q Fi

842_ ܂ simultaneously, two BR Solving B. ܕܪ- ao . ܀ . ܕ.- £[a-:] (.܀ ܨܪܐ) - ܘ * ܆ ܀ $ 4, - «. - 30 + 4 ܘ: - ; - ܙ: ܀ _ .: ; ܙ